# Infrared Dressing and the Strong CP Problem: Geometric Renormalization of the Vacuum Angle

### v1

@@ I Introduction
In non-Abelian gauge theories, the parameter $\theta$ leaves the local Yang–Mills equations unchanged but affects how gauge configurations belonging to different homotopy classes are coherently combined in the quantum ground state. The observed suppression of strong CP violation may therefore be interpreted as an indication that the infrared sector selects a vacuum representation in which the global phase associated with these topological sectors becomes dynamically ineffective Belavin1975 ; JackiwRebbi1976 ; tHooft1976 ; Witten1979 ; Veneziano:1979ec ; PecceiQuinn1977 .
Within the functional integral, this parameter enters as a weight assigned to inequivalent winding sectors. However, such a parametrization does not reflect the reorganization of physical states that occurs once slow collective motion between topological configurations is separated from fast gluonic fluctuations in the infrared regime, as emphasized in the infrared dressing framework developed in Gamboa:2025hxa .
After implementing this adiabatic separation, the reduced dynamics is governed by an effective Hamiltonian defined along a compact collective coordinate related to the Chern–Simons number. In general, the quantity that controls this reduced infrared evolution differs from the bare vacuum angle, as it includes a geometric contribution induced by the response of the fast gluonic sector.
In this framework, the apparent insensitivity of certain local probes to $\theta$ does not imply a trivial vacuum structure, but instead reflects the fact that such observables act within a fixed infrared representation. Only operators sensitive to global boundary data can detect the corresponding holonomy that characterizes inequivalent infrared sectors.
@@ V Quantitative Benchmark: $\theta$-Dependence, Order of Limits, and the Role of Quarks
$\chi_{\rm top}=\left.\frac{\partial^{2}f(\theta)}{\partial\theta^{2}}\right|_{\theta=0}=\frac{1}{V},\quad\langle Q^{2}\rangle_{\theta=0}=\int d^{4}x\,\langle q(x)\,q(0)\rangle_{\theta=0},\qquad q(x)=\frac{1}{32\pi^{2}}\,\mathrm{tr}\,F\widetilde{F}$. (21)
@@
A sharp benchmark illustrating this distinction is provided by the quantum rotor toy model, where $\theta$ is unambiguously realized as a holonomy through a twisted boundary condition. In this setting, both the vacuum energy and the topological susceptibility can be reconstructed nonperturbatively from local correlation functions AlbandeaCatumbaRamosPRD , confirming that global response functions remain sensitive to the holonomy even when many local observables are blind to it. (see also, Benabou:2025viy )
@@ VIII Induced Berry holonomy along the Chern–Simons collective coordinate
The analysis of the previous section identifies the physically relevant parameter governing the infrared vacuum structure as the effective holonomy $\theta_{\rm eff}$ defined in Eq. (40). However, at this stage $\theta_{\rm eff}$ is introduced at the level of the reduced Born–Oppenheimer Hamiltonian and its dynamical origin remains to be established. In particular, it is not a priori clear whether the additional term appearing in Eq. (40) represents a mere reparametrization of the vacuum angle or a genuine infrared contribution induced by the dressing of fast gluonic modes.
This distinction is crucial from the viewpoint of the strong CP problem. If the Born–Oppenheimer connection entering the reduced Hamiltonian were trivial along the compact collective coordinate associated with topological transitions, the effective parameter $\theta_{\rm eff}$ would simply reproduce the bare vacuum angle $\theta$, and no dynamical mechanism would emerge from the infrared sector. Conversely, a non–trivial Berry holonomy acquired by the fast gluonic modes during adiabatic transport between topologically inequivalent configurations would induce a shift of the physical holonomy that enters the infrared Hamiltonian.
The purpose of this section is to show that, in Yang–Mills theory, the Born–Oppenheimer connection associated with the Chern–Simons collective coordinate is generically non–vanishing. As a consequence, adiabatic transport of the fast sector along a closed contour in configuration space induces a Berry phase that modifies the physical vacuum angle appearing in the infrared dynamics. In this sense, the vacuum angle becomes geometrically renormalized by infrared dressing, and the strong CP problem may be reformulated as a question of vacuum selection governed by the dynamically induced effective holonomy $\theta_{\rm eff}$.
In order to assess whether the effective holonomy defined in Eq. (28) represents a purely formal reparametrization or a dynamically induced quantity, one must determine whether the induced Born–Oppenheimer connection
@@
This parametrization is not an additional dynamical assumption, but reflects the possible topological obstruction to defining a globally single–valued fast–sector ground state over the reduced configuration space.
In the adiabatic construction, the fast gluonic vacuum defines a line bundle over the space of slow collective configurations. If this bundle fails to admit a global section along the Chern–Simons cycle—which parametrizes motion between topologically inequivalent vacua—then adiabatic transport around that cycle necessarily produces a non–trivial Berry phase.
Equivalently, a non–vanishing pullback of the functional Berry curvature onto the one–cycle generated by the Chern–Simons collective coordinate leads to a holonomy proportional to the winding number $Q$. In this case, the integral of the induced connection along the cycle cannot be removed by any admissible local redefinition of the fast–sector vacuum, and therefore defines a genuine infrared observable.
Therefore,
$\oint d\phi\;A_{\rm BO}(\phi)=\kappa\,Q,$ (41)
Thus the induced Berry holonomy produces a shift of the physical vacuum parameter entering the infrared Hamiltonian, summarized by the effective holonomy $\theta_{\rm eff}$ in Eq. (28).
In this sense, the vacuum angle is geometrically renormalized by infrared dressing: adiabatic transport of the fast gluonic sector along a topologically non–trivial contour in configuration space induces a Berry phase that shifts the physical holonomy.
IX Self-consistent infrared relaxation of the effective holonomy
Equations (39) and (40) show that infrared vacuum selection is governed by the effective holonomy $\theta_{\rm eff}$ rather than by the bare parameter $\theta$. However, for this statement to amount to a dynamical solution of the strong CP problem, the Berry contribution cannot be treated as an arbitrary shift. Instead, it must be determined self–consistently by the fast gluonic sector that has been integrated out.
In the Born–Oppenheimer framework, the induced connection $A_{\rm BO}(\phi)=i\langle n(\phi)|\partial_{\phi}|n(\phi)\rangle$ is not an external input: it is a functional of the fast–sector ground state computed in the presence of the physical holonomy entering the reduced Hamiltonian. Since the fast eigenproblem is defined with the same global data that controls infrared boundary conditions, the induced holonomy must be regarded as a response function,
@@ IX Self-consistent infrared relaxation of the effective holonomy
so that the definition of $\theta_{\rm eff}$ becomes a fixed–point equation,
@@
The mechanism is transparent at the level of vacuum selection. The total ground–state energy may be organized as
$E_{\rm tot}(\theta_{\rm eff})=E_{0}(\theta_{\rm eff})+E_{\rm fast}(\theta_{\rm eff}),$ (45)
where $E_{0}$ denotes the infrared rotor contribution and $E_{\rm fast}$ is the fast–sector contribution generated by integrating out non–collective gluonic modes. Since $\theta_{\rm eff}$ depends on $\kappa$, minimizing the vacuum energy with respect to the fast–sector dressing yields the self–consistency condition
$0=\frac{\partial E_{\rm tot}}{\partial\kappa}=(2\pi Q)\,\frac{\partial E_{\rm tot}}{\partial\theta_{\rm eff}}+\left.\frac{\partial E_{\rm fast}}{\partial\kappa}\right|_{\theta_{\rm eff}},$ (46)
[…]
X Infrared Renormalization Group Flow of the Effective Vacuum Angle
The self–consistency condition obtained in the previous section implies that the effective holonomy $\theta_{\rm eff}$ is determined by the response of the fast gluonic sector through the functional relation
$\theta_{\rm eff}=\theta+2\pi\,Q\,\kappa(\theta_{\rm eff}),$ (47)
where $\kappa$ encodes the Berry response induced by infrared dressing.
This relation may be interpreted dynamically by introducing an infrared scale parameter $s=\ln\mu$ associated with the adiabatic separation between slow topological modes and fast gluonic fluctuations. In this framework, the induced dressing of the fast sector depends on the scale at which the separation is implemented, so that the effective holonomy becomes a scale–dependent quantity,
$\theta_{\rm eff}\;\longrightarrow\;\theta_{\rm eff}(s).$ (48)
In order to obtain a dynamical infrared flow, one must take into account that the induced response of the fast sector depends on the scale at which the Born–Oppenheimer separation is implemented. Accordingly, the Berry response becomes a scale–dependent quantity,
@@ X Infrared Renormalization Group Flow of the Effective Vacuum Angle
$\frac{d\theta_{\rm eff}}{ds}=2\pi Q\left[\frac{\partial\kappa}{\partial\theta_{\rm eff}}\frac{d\theta_{\rm eff}}{ds}+\frac{\partial\kappa}{\partial s}\right],$ (51)
which may be solved for the infrared beta function,
$\beta(\theta_{\rm eff})\equiv\frac{d\theta_{\rm eff}}{ds}=\frac{2\pi Q\,\partial_{s}\kappa(\theta_{\rm eff},s)}{1-2\pi Q\,\partial_{\theta_{\rm eff}}\kappa(\theta_{\rm eff},s)}.$ (52)
[…]
Linearizing around $\theta_{\rm eff}^{\star}$ with $\delta\theta_{\rm eff}:=\theta_{\rm eff}-\theta_{\rm eff}^{\star}$, one finds
$\frac{d\,\delta\theta_{\rm eff}}{ds}=\beta^{\prime}(\theta_{\rm eff}^{\star})\,\delta\theta_{\rm eff}\qquad\Rightarrow\qquad\delta\theta_{\rm eff}(s)\propto e^{\,\beta^{\prime}(\theta_{\rm eff}^{\star})s}.$ (53)
Since the infrared corresponds to $s\to$-$\infty$, the fixed point is infrared attractive if
$\beta^{\prime}(\theta_{\rm eff}^{\star})>0,$ (54)
so that $\delta\theta_{\rm eff}(s)\to 0$ along the infrared flow.
[…]
$\beta^{\prime}(\theta_{\rm eff})=\frac{2\pi Q\,\kappa^{\prime\prime}(\theta_{\rm eff})}{\bigl(1-2\pi Q\,\kappa^{\prime}(\theta_{\rm eff})\bigr)^{2}}.$ (55)
At a fixed point where $\kappa^{\prime}(\theta_{\rm eff}^{\star})=0$, this reduces to $\beta^{\prime}(\theta_{\rm eff}^{\star})=2\pi Q\,\kappa^{\prime\prime}(\theta_{\rm eff}^{\star})$. Therefore the CP–invariant fixed point $\theta_{\rm eff}^{\star}=0$ (mod $2\pi$) is infrared attractive whenever the induced response satisfies $Q\,\kappa^{\prime\prime}(0)>0$.
A CP–invariant configuration corresponds to a fixed point of the infrared flow,
$\beta(\theta_{\rm eff}^{\star})=0,$ (56)
which occurs whenever
$\kappa^{\prime}(\theta_{\rm eff}^{\star})=0.$ (57)
If the infrared response of the fast gluonic sector admits an isolated extremum at $\theta_{\rm eff}=0$ (mod $2\pi$), the induced renormalization group flow drives the effective holonomy toward a CP–invariant value in the infrared limit,
$\theta_{\rm eff}(s)\;\longrightarrow\;0,\qquad s\to-\infty.$ (58)
In this sense, CP violation is not forbidden by the fundamental theory but becomes dynamically suppressed along the infrared renormalization group flow generated by adiabatic dressing. The strong CP problem may therefore be reformulated as a nonperturbative infrared relaxation mechanism governed by the Berry response of the fast gluonic sector.
XI Conclusion
In this work we have reexamined the strong CP problem from the viewpoint of the infrared structure of Yang–Mills theory. Rather than treating the vacuum angle $\theta$ as a fixed external parameter entering the functional integral, we have shown that the physically relevant quantity governing the infrared vacuum dynamics is the effective holonomy $\theta_{\rm eff}$ defined in Eq. (28), which appears in the reduced Born–Oppenheimer Hamiltonian describing slow motion along the compact collective coordinate associated with topological transitions.
As shown in Section VIII, the Berry connection induced by the fast gluonic sector along the Chern–Simons collective coordinate is generically non–trivial. As a result, adiabatic transport of the fast modes along closed contours in configuration space induces a Berry phase that shifts the physical holonomy entering the infrared Hamiltonian. In this sense, the vacuum angle becomes geometrically renormalized by infrared dressing.
The self–consistency condition relating the induced Berry response to the effective holonomy defines a nonperturbative infrared renormalization group flow for $\theta_{\rm eff}$. As shown in Section IX, this flow admits CP–invariant fixed points characterized by

### v2

# arXiv:2603.00220v2[hep-th] 03 Mar 2026

@@ I Introduction
In non–Abelian gauge theories, CP violation is not associated with local processes but with the global structure of the quantum vacuum Belavin1975 ; JackiwRebbi1976 ; tHooft1976 ; Witten1979 ; Veneziano:1979ec ; PecceiQuinn1977 . The vacuum angle $\theta$ does not modify the classical Yang–Mills equations, but controls the relative phase with which gauge configurations belonging to distinct topological sectors are combined in the quantum ground state. The observed suppression of strong CP violation therefore raises a question of vacuum selection rather than of local dynamics.
In the standard functional integral formulation, the parameter $\theta$ enters as a weight assigned to inequivalent winding sectors. This parametrization, however, does not reflect the reorganization of physical states that occurs in the infrared regime, where slow collective motion between topological configurations becomes separated from fast gluonic fluctuations. As emphasized in the infrared dressing framework developed in Gamboa:2025hxa , such an adiabatic separation leads to a reduced Born–Oppenheimer dynamics along a compact collective coordinate related to the Chern–Simons number.
In general, the quantity that controls this reduced infrared evolution differs from the bare vacuum angle, as it includes a geometric contribution induced by the response of the fast gluonic sector. In this framework, the apparent insensitivity of certain local probes to $\theta$ does not imply a trivial vacuum structure, but instead reflects the fact that such observables act within a fixed infrared representation. Only operators sensitive to global boundary data can detect the holonomy that characterizes inequivalent infrared sectors.
@@
In what follows, we develop this perspective in a sequence of steps. We first identify the topological collective coordinate governing slow infrared motion, then implement an adiabatic separation between collective and gluonic degrees of freedom. The geometric response of the fast sector induces a Berry connection along this coordinate, leading to a self–consistent effective vacuum angle $\theta_{\rm eff}$. Finally, we show that this consistency condition may be interpreted as an infrared renormalization group flow which dynamically suppresses CP violation.
@@ V Quantitative Benchmark: $\theta$-Dependence, Order of Limits, and the Role of Quarks
$\chi_{\rm top}=\left.\frac{\partial^{2}f(\theta)}{\partial\theta^{2}}\right|_{\theta=0}=\frac{1}{V}\,\langle Q^{2}\rangle_{\theta=0}=\int d^{4}x\,\langle q(x)\,q(0)\rangle_{\theta=0},\qquad q(x)=\frac{1}{32\pi^{2}}\,\mathrm{tr}\,F\widetilde{F}$. (21)
@@
A sharp benchmark illustrating this distinction is provided by the quantum rotor toy model, where $\theta$ is unambiguously realized as a holonomy through a twisted boundary condition. In this setting, both the vacuum energy and the topological susceptibility can be reconstructed nonperturbatively from local correlation functions AlbandeaCatumbaRamosPRD , confirming that global response functions remain sensitive to the holonomy even when many local observables are blind to it (see also Benabou:2025viy ).
@@ VIII Induced Berry holonomy along the Chern–Simons collective coordinate
[…]
This distinction is crucial from the viewpoint of the strong CP problem. If the induced connection were trivial along the Chern–Simons cycle, then $\theta_{\rm eff}$ would reduce to $\theta$ and no dynamical mechanism would emerge from the infrared sector. Conversely, a non–trivial Berry holonomy acquired by the fast gluonic modes during adiabatic transport between topologically inequivalent configurations induces a shift of the physical holonomy entering the infrared Hamiltonian.
In order to assess this, one must determine whether the induced Born–Oppenheimer connection
@@
Equivalently, a non–vanishing pullback of the functional Berry curvature onto the one–cycle generated by the Chern–Simons collective coordinate leads to a holonomy proportional to the winding number $Q$. In this case, the integral of the induced connection along the cycle cannot be removed by any admissible local redefinition of the fast–sector vacuum, and therefore defines a genuine infrared observable:
$\oint d\phi\;A_{\rm BO}(\phi)=\kappa\,Q.$ (41)
[…]
IX Infrared Relaxation and Renormalization of the Effective Holonomy
The emergence of an effective holonomy $\theta_{\rm eff}$ makes vacuum selection a genuinely infrared dynamical question. For this to amount to a mechanism for strong CP, the Berry contribution cannot be treated as an arbitrary shift: it must be determined self–consistently by the fast gluonic sector that has been integrated out.
In the Born–Oppenheimer framework, the induced holonomy depends on the fast–sector ground state computed in the presence of the physical holonomy entering the reduced Hamiltonian. The induced Berry holonomy must therefore be regarded as a response function,
@@ IX Infrared Relaxation and Renormalization of the Effective Holonomy
so that the definition of $\theta_{\rm eff}$ becomes a fixed–point condition,
@@
This relaxation mechanism admits a dynamical interpretation by introducing an infrared scale parameter $s=\ln\mu$ associated with the adiabatic separation between slow topological modes and fast gluonic fluctuations. The induced dressing depends on the scale at which the separation is implemented, so that the response becomes scale dependent,
@@
Differentiating with respect to $s$ yields an induced infrared flow equation,
$\frac{d\theta_{\rm eff}}{ds}=\frac{2\pi Q\,\partial_{s}\kappa(\theta_{\rm eff},s)}{1-2\pi Q\,\partial_{\theta_{\rm eff}}\kappa(\theta_{\rm eff},s)}.$ (47)
In the present convention $s=\ln\mu$, infrared evolution corresponds to $s\to$-$\infty$. Linearizing around a fixed point $\theta_{\rm eff}^{\star}$ with $\delta\theta_{\rm eff}:=\theta_{\rm eff}-\theta_{\rm eff}^{\star}$, one finds
$\delta\theta_{\rm eff}(s)\propto e^{\,\beta^{\prime}(\theta_{\rm eff}^{\star})s}.$ (48)
The fixed point is therefore infrared attractive whenever
$\beta^{\prime}(\theta_{\rm eff}^{\star})>0.$ (49)
At a CP–invariant extremum satisfying $\partial_{\theta_{\rm eff}}\kappa(\theta_{\rm eff}^{\star})=0$, this condition reduces to
$Q\,\kappa^{\prime\prime}(\theta_{\rm eff}^{\star})>0,$ (50)
so that the infrared flow drives
$\theta_{\rm eff}(s)\;\longrightarrow\;0,\qquad s\to-\infty.$ (51)
In this sense, CP violation is not forbidden by the microscopic theory but becomes dynamically suppressed along the infrared renormalization group flow generated by adiabatic dressing. The strong CP problem may therefore be reformulated as a nonperturbative infrared relaxation mechanism governed by the Berry response of the fast gluonic sector.
X Conclusion
@@
As shown in Sec. VIII, the Berry connection induced by the fast gluonic sector along the Chern–Simons collective coordinate is generically non–trivial. As a result, adiabatic transport of the fast modes along closed contours in configuration space induces a Berry phase that shifts the physical holonomy entering the infrared Hamiltonian. In this sense, the vacuum angle becomes geometrically renormalized by infrared dressing.
The self–consistency condition relating the induced Berry response to the effective holonomy defines a nonperturbative infrared renormalization group flow for $\theta_{\rm eff}$. As shown in Sec. IX, this flow admits CP–invariant fixed points characterized by